### v1

# Samer Al-Hamadani

Abstract
Object detection systems have traditionally relied on supervised learning with manually annotated bounding boxes, achieving high accuracy at the cost of substantial annotation investment. The emergence of Vision-Language Models (VLMs) offers an alternative paradigm enabling zero-shot detection through natural language queries, eliminating annotation requirements but operating with reduced accuracy. This paper presents the first comprehensive cost-effectiveness analysis comparing supervised detection (YOLO) with zero-shot VLM inference (Gemini Flash 2.5). Through systematic evaluation on 1,000 stratified COCO images and 200 diverse product images spanning consumer electronics and rare categories, combined with detailed Total Cost of Ownership modeling, we establish quantitative break-even thresholds governing architecture selection. Our findings reveal that supervised YOLO achieves 91.2% accuracy versus 68.5% for zero-shot Gemini on standard categories, representing a 22.7 percentage point advantage that costs \$10,800 in annotation for 100-category systems. However, this advantage justifies investment only beyond 55 million inferences, equivalent to 151,000 images daily for one year. Zero-shot Gemini demonstrates 52.3% accuracy on diverse product categories (ranging from highly web-prevalent consumer electronics at 75–85% to rare specialized equipment at 25–40%) where supervised YOLO achieves 0% due to architectural constraints preventing detection of untrained classes. Cost per Correct Detection analysis reveals substantially lower per-detection costs for Gemini (\$0.00050 vs \$0.143) at 100,000 inferences despite accuracy deficits. We develop decision frameworks demonstrating that optimal architecture selection depends critically on deployment volume, category stability, budget constraints, and accuracy requirements rather than purely technical performance metrics.
Keywords: Cost-Effectiveness Analysis, Object Detection, Total Cost of Ownership, Vision-Language Models, Zero-Shot Learning, YOLO
1 Introduction
Object detection constitutes a foundational computer vision capability enabling diverse applications from autonomous vehicles to retail analytics, with modern deep learning approaches achieving remarkable technical performance exceeding 90% mean Average Precision on standardized benchmarks [1, 2]. However, technical accuracy represents only one dimension of deployment viability, as real-world system selection requires evaluating cost-effectiveness—the relationship between detection performance and total economic investment required to achieve that performance [3, 4]. Traditional supervised detectors, exemplified by the YOLO architecture family [2, 5], rely fundamentally on manually annotated training data, with industry reports estimating annotation costs between \$0.10 and \$0.50 per bounding box [6, 7], translating to \$9,000–\$45,000 for establishing 100-category detection systems with sufficient training data.
Vision-Language Models represent an alternative paradigm achieving object detection through zero-shot inference without task-specific supervision [8, 9, 10]. Pre-trained on billions of image-text pairs, VLMs accept natural language object descriptions and generate bounding box predictions through learned visual-linguistic alignment, fundamentally eliminating annotation requirements. Commercial VLM APIs from providers including Google Gemini, OpenAI GPT-4V, and Anthropic Claude offer pay-per-use pricing models with per-image costs as low as \$0.00025, representing a fundamentally different economic structure compared to supervised approaches requiring substantial upfront annotation investment. The contrast between these paradigms raises critical questions about practical deployment: when does the accuracy advantage of supervised training justify its annotation costs? How do zero-shot VLMs perform on categories completely absent from training data? What deployment scenarios favor each approach from a cost-benefit perspective?
Despite the emergence of these two fundamentally different detection paradigms with distinct cost structures and performance characteristics, systematic cost-effectiveness analysis remains absent from existing literature. Current research focuses predominantly on accuracy benchmarking [1, 11], with economic considerations relegated to brief discussion sections if addressed at all. Critical questions persist regarding the practical deployment conditions under which each paradigm achieves optimal return on investment. At what inference volume do annotation costs justify supervised training investment? How do break-even thresholds vary across deployment scenarios with different budget constraints, accuracy requirements, and category evolution rates? Which paradigm optimizes cost-effectiveness for organizations operating under specific operational constraints?
This paper addresses these gaps through comprehensive cost-effectiveness analysis comparing supervised YOLO with zero-shot Gemini VLM across technical performance and economic dimensions. We develop detailed Total Cost of Ownership models quantifying annotation expenses, training infrastructure costs, inference pricing, and maintenance overhead grounded in industry pricing data. Through experiments on 1,000 stratified COCO images (20% of validation set) and 200 diverse product images spanning highly web-prevalent consumer electronics to specialized rare equipment, we establish quantitative thresholds where each paradigm achieves cost-effectiveness, demonstrating that zero-shot VLMs maintain economic advantages below 55 million images while supervised approaches dominate beyond this threshold. Our expanded evaluation provides stronger statistical power while acknowledging that full 5,000-image COCO validation would further narrow confidence intervals. Our analysis of representative deployment contexts provides architecture selection guidelines based on deployment-specific constraints, enabling practitioners to select architectures through quantitative break-even analysis rather than subjective performance assessments.
Figure 1: Architectural comparison of supervised YOLO versus zero-shot Gemini VLM.
2 Related Work
The YOLO architecture introduced by Redmon et al. [2] reformulated object detection as unified regression, achieving real-time performance through single-stage processing. Subsequent iterations refined this foundation through architectural improvements including multi-scale predictions in YOLOv3 [12], CSPDarknet backbone in YOLOv4 [13], and anchor-free detection in YOLOv8 [5]. These supervised detectors require extensive annotated training data, with the COCO dataset [1] providing 118,287 images containing 860,001 bounding box annotations across 80 categories as the standard training corpus. The fundamental training objective optimizes a composite loss function balancing localization accuracy, objectness confidence, and classification correctness [14]:
$L_{\text{total}}=\lambda_{\text{coord}}L_{\text{coord}}+\lambda_{\text{obj}}L_{\text{obj}}+\lambda_{\text{cls}}L_{\text{cls}}$ (1)
where $L_{\text{coord}}$ measures localization error, $L_{\text{obj}}$ penalizes confidence errors, $L_{\text{cls}}$ captures classification errors, and $\lambda_{\text{coord}}$, $\lambda_{\text{obj}}$, $\lambda_{\text{cls}}$ are the hyperparameters controlling the contribution of each term. Coordinate loss measures bounding box accuracy, objectness loss assesses detection confidence, and classification loss evaluates category prediction correctness. This formulation critically requires ground-truth bounding box annotations for every training instance, with annotation costs representing the primary economic barrier to supervised detection deployment [15].
Vision-Language Models pioneered by CLIP [8] demonstrated that training on 400 million image-text pairs enables transferable visual representations through contrastive learning objectives aligning visual and linguistic embeddings:
$L_{\text{contrastive}}=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp(\text{sim}(I_{i},T_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(I_{i},T_{j})/\tau)}\right]$ (2)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity between image embedding $I$ and text embedding $T$, with temperature parameter $\tau$ controlling distribution sharpness. This objective encourages correct image-text pairings to achieve higher similarity than incorrect combinations, creating shared embedding spaces enabling zero-shot transfer. Subsequent architectures extended these foundations, with Flamingo [9] interleaving vision encoders with autoregressive language models, and Gemini [10, 16] employing multimodal-native training where vision and language components undergo joint optimization from initialization. These VLMs demonstrate object localization through bounding box coordinate generation conditioned on natural language queries, enabling zero-shot detection without task-specific training on annotated detection datasets.
The zero-shot capability emerges from training on diverse image-text pairs where natural language captions frequently contain spatial expressions like “the car on the left side” or “person wearing blue shirt,” providing implicit supervision for spatial grounding without explicit bounding box annotations. When presented with novel categories like “AirPods Pro” during inference, VLMs retrieve learned visual concepts associated with “wireless earbuds,” “charging case,” and “Apple design” through linguistic similarity, enabling localization despite never encountering explicit “AirPods Pro” training examples with bounding boxes. This architectural flexibility contrasts fundamentally with supervised detectors constrained to fixed category indices defined during training.
Sculley et al. [3] highlighted that machine learning system costs extend far beyond model training, with data collection, infrastructure management, and maintenance comprising over 90% of total investment in production environments. Sambasivan et al. [15] documented annotation challenges through practitioner interviews, revealing that data quality issues and labeling overhead constitute primary bottlenecks, with annotation costs frequently exceeding model development expenses particularly for specialized domains requiring expert labeling. Their findings indicate that in medical imaging contexts, expert radiologist annotation at \$150–\$300 per hour compared to general crowdsourced annotation at \$15–\$30 per hour drives per-box costs to \$1.00 or higher, making supervised approaches economically prohibitive for many organizations. Industry reports provide concrete pricing benchmarks establishing the economic landscape for detection deployment. Scale AI [7] documents annotation costs ranging from \$0.10 per bounding box for simple objects with clear boundaries to \$0.50 per box for complex scenes requiring multiple annotators with consensus resolution and quality verification. Appen [6] reports similar ranges with additional platform fees averaging 5% of base annotation costs. Cloud GPU pricing from major providers establishes training infrastructure costs, with NVIDIA A100 instances averaging \$1.50 to \$3.00 per hour across AWS, Google Cloud, and Azure as of October 2024 [17]. Commercial VLM API pricing varies substantially across providers, with Gemini Flash charging \$0.00025 per image, GPT-4V costing \$0.01 per image representing a 40-fold premium, and Claude 3 pricing at \$0.008 per image [18, 19, 20].
Despite extensive research on detection accuracy and growing recognition of deployment costs, no prior work systematically analyzes cost-effectiveness trade-offs between supervised and zero-shot detection paradigms. Existing benchmarks report mean Average Precision and inference latency without consideration of annotation investments or operational expenses [1]. Studies comparing supervised and zero-shot approaches focus on accuracy differentials without quantifying the break-even volumes where annotation investments prove economically justifiable [8]. Our work addresses this gap by combining technical evaluation with comprehensive Total Cost of Ownership modeling, enabling quantitative architecture selection based on deployment-specific constraints rather than purely technical performance metrics.
3 Cost Modeling Framework
We develop comprehensive Total Cost of Ownership models quantifying expenses across both detection paradigms, categorizing costs into annotation, training, infrastructure, and inference components. All cost estimates derive from publicly available industry pricing data with currency in US dollars. Our framework enables practitioners to calculate expected costs for specific deployment scenarios by substituting organization-specific parameters including category counts, daily inference volumes, and accuracy requirements.
For supervised detection using YOLO architecture, annotation costs represent the primary expense barrier. Following industry pricing analysis [7, 6], we model total annotation cost as:
$C_{\text{annotation}}=N_{\text{cat}}\times N_{\text{img/cat}}\times N_{\text{box/img}}\times P_{\text{box}}\times 1.20$ (3)
where $N_{\text{cat}}$ denotes category count, $N_{\text{img/cat}}$ represents images per category (typically 100–500 for deep learning), $N_{\text{box/img}}$ averages boxes per image (approximately 3 for focused detection), $P_{\text{box}}$ equals price per box (\$0.30 baseline for standard quality), and the 1.20 factor accounts for 15% quality control plus 5% platform fees. For a 100-category system with 100 images per category, total annotation investment reaches \$10,800. Training costs encompass GPU compute at approximately \$2 per hour for 8 hours plus \$300 engineering overhead for hyperparameter tuning and validation, totaling \$316 for enterprise deployments. Infrastructure costs for edge deployment using devices like NVIDIA Jetson Orin (\$500) amortize over three-year lifetime and daily inference volume, calculating per-image cost as hardware price divided by total lifetime inferences.
The supervised detection TCO equation becomes:
$\text{TCO}_{\text{YOLO}}(N)=C_{\text{annotation}}+C_{\text{training}}+C_{\text{infrastructure}}+N\times C_{\text{inference}}$ (4)
where $N$ represents total inference count and $C_{\text{inference}}$ reflects amortized infrastructure costs typically under \$0.0001 per image for high-volume deployments after hardware amortization.
Zero-shot VLMs operate through API inference pricing without upfront costs, with Gemini Flash 2.5 charging \$0.00025 per image:
$\text{TCO}_{\text{Gemini}}(N)=N\times C_{\text{API}}$ (5)
where $C_{\text{API}}=\mathdollar 0.00025$ represents per-image API pricing. Gemini provides generous free-tier access allowing 1,500 requests daily, enabling zero-cost prototyping for applications processing under this threshold.
We derive break-even inference volume $N^{*}$ where supervised and VLM TCO equalize by setting equations equal:
$C_{\text{annotation}}+C_{\text{training}}+C_{\text{infrastructure}}=N^{*}\times(C_{\text{API}}-C_{\text{inference}})$ (6)
Solving for 100-category system with \$10,800 annotation, \$316 training, \$500 infrastructure, negligible supervised inference cost after amortization, and \$0.00025 VLM cost:
$N^{*}=\frac{11,616}{0.00025-0}\approx 46,464,000\text{ images}$ (7)
However, accounting for realistic infrastructure amortization at high volumes where supervised per-image cost reaches approximately \$0.00004 (from hardware replacement and maintenance), break-even adjusts to:
$N^{*}=\frac{11,616}{0.00025-0.00004}=55,314,286\text{ images}$ (8)
This translates to processing 151,500 images daily for one year, or 15,150 images daily sustained over ten years, revealing that annotation investments justify costs only for extreme-volume production systems.
Important Note on Break-Even Stability: Our break-even calculations assume stable pricing structures over deployment lifetimes. However, commercial VLM API pricing exhibits high volatility, with Gemini Flash costs declining 80% from \$0.00125 (October 2023) to \$0.00025 (October 2024) within one year. Continued price reductions would shift break-even thresholds to progressively higher volumes, potentially extending VLM cost-effectiveness regions. Conversely, annotation costs have remained relatively stable over the past decade, with crowdsourced platforms maintaining \$0.10–\$0.50 per box ranges. Organizations should recalculate break-even thresholds annually as the VLM pricing landscape evolves, rather than treating our calculated 55 million image threshold as fixed.
To integrate cost and performance, we define Cost per Correct Detection:
$\text{CCD}=\frac{\text{TCO}(N)}{N\times\text{Accuracy@IoU}_{0.5}}$ (9)
This metric penalizes both high costs and low accuracy, enabling fair comparison between supervised approaches with high upfront costs and high accuracy versus zero-shot approaches with low operational costs and moderate accuracy. Lower CCD indicates superior cost-effectiveness by quantifying dollars spent per successfully detected object.
Table 1 summarizes cost parameters and break-even analysis across system scales, illustrating how annotation investment scales with category count while dramatically affecting economic viability.
Table 1: Annotation cost structure and break-even performance across different system scales
System Categories Images Annotation Training Total Break-Even
Scale +Infra Upfront Vol / Daily (1yr)
Small 10 1,000 \$1,080 \$554 \$1,634 7.8M / 21,300
Medium 50 5,000 \$5,400 \$680 \$6,080 28.9M / 79,200
Large 100 10,000 \$10,800 \$816 \$11,616 55.3M / 151,500
Enterprise 200 30,000 \$21,600 \$1,000 \$22,600 107.6M / 294,800
Medical 100 10,000 \$36,000 \$816 \$36,816 175.3M / 480,300
Specifically, it details the cost structure and break-even points under different assumptions, where the medical domain assumes \$1.00/box for expert radiologist annotation versus \$0.30/box standard. The break-even volume represents the inference count at which YOLO and Gemini total costs equalize, while the daily volume column indicates the number of images per day required to reach break-even within one year. Notably, most real-world applications operate far below these thresholds, economically favoring the zero-shot Gemini.
Figure 2 visualizes TCO evolution across inference volumes for 100-category systems, illustrating fundamental differences in cost structures between upfront-heavy supervised training and pay-per-use zero-shot inference.
Figure 2: Total Cost of Ownership evolution across inference volumes for 100-category detection system.
4 Methodology
Our experimental design evaluates technical performance across both detection paradigms through systematic testing on standard and novel object categories. We construct a stratified 1,000-image subset from COCO 2017 validation set using random sampling preserving object size distributions defined by Lin et al [1]. Objects categorize as small (area << 1,024 pixels), medium (1,024–9,216 pixels), or large (>> 9,216 pixels), with sampling targeting 200 small-object images (20%), 400 medium (40%), and 400 large (40%) matching full COCO proportions. Random seed 42 ensures reproducibility, enabling reconstruction of our exact evaluation subset from public COCO dataset. This 20% COCO sample size provides 80% statistical power to detect accuracy differences $\geq$3 percentage points at $\alpha=0.05$. This stratification ensures performance measurements reflect capabilities across diverse object scales rather than biasing toward predominantly large objects that prove easier to detect.
For novel category evaluation testing genuine zero-shot generalization, we curate 200 images spanning 20 diverse product categories intentionally selected to represent spectrum from highly web-prevalent consumer electronics to specialized rare equipment. Our selection strategy includes:
Tier 1 - Highly Web-Prevalent (10 categories, 10 images each): post-2017 consumer products with millions of online images including AirPods Pro wireless earbuds (released October 2019), Amazon Echo Dot 4th generation smart speaker (2020), Ring Video Doorbell, Tesla Model 3 electric vehicle (July 2017), Peloton connected fitness bikes, Google Nest Hub smart displays (2019), Oculus Quest 2 VR headsets (October 2020), DJI Mavic drones, Nintendo Switch OLED, and Apple AirTag trackers. These products benefit from extensive product marketing, user-generated content, unboxing videos, and e-commerce imagery.
Tier 2 - Moderately Web-Prevalent (5 categories, 10 images each): Products with moderate online presence including electric scooters, N95 respirator masks, height-adjustable standing desks, smart door locks, and portable power stations. These categories have thousands rather than millions of online images.
Tier 3 - Specialized/Rare (5 categories, 10 images each): Equipment with minimal web representation including industrial 3D metal printers (specialized manufacturing), underwater ROV inspection equipment, precision laboratory pipettes, professional broadcast cameras (specialized models), and medical diagnostic ultrasound transducers. These categories intentionally test true zero-shot capability on objects with fewer than 500 online images, reflecting genuine deployment scenarios for specialized domains.
Images source from Creative Commons repositories (Unsplash, Pexels, Wikimedia), industry databases, and specialized equipment catalogs with verified licensing. Three expert annotators consisting of the paper authors plus two domain specialists for Tier 3 equipment independently draw ground truth bounding boxes using LabelImg annotation software [21]. Annotation instructions specify tight bounding boxes encompassing complete objects with minimal background inclusion. Inter-annotator agreement measured via Fleiss’ kappa reaches 0.89 for Tier 1, 0.84 for Tier 2, and 0.78 for Tier 3, with lower agreement on specialized equipment reflecting genuine annotation difficulty for rare objects.
Model configurations employ YOLOv8m as supervised baseline, utilizing the medium variant with 25.9 million parameters pre-trained on COCO train2017 containing 118,287 images. Implementation uses Ultralytics YOLOv8 library version 8.0.196 with confidence threshold 0.25, IoU threshold 0.7 for non-maximum suppression, $\times640$ input size, and FP16 mixed precision on NVIDIA T4 GPU accessed through Google Colaboratory Pro. Evaluation completes in approximately 92 minutes for 1,000 COCO images at batch size 16.
For zero-shot evaluation, Gemini Flash 2.5 accesses through Google AI Studio API with temperature 0.1 for deterministic outputs, maximum tokens 512, employing chain-of-thought prompting optimized through preliminary experiments following Wei et al. [22]. Prompt structure requests step-by-step reasoning before JSON output containing bounding box coordinates verified within image dimensions. API evaluation requires 1,200 total requests (1,000 COCO + 200 novel) distributed across 8 days (October 1–9, 2025) using free-tier quotas (15 requests/minute, 1,500/day) across three researcher accounts, staying within rate limits while maintaining consistent model version.
Evaluation metrics include Intersection over Union quantifying bounding box overlap [23], accuracy at thresholds 0.5 and 0.7 following COCO conventions, mean IoU averaging overlap across predictions, and inference latency measuring wall-clock time. Statistical analysis employs paired t-tests ($\alpha=0.05$) for significance testing, Cohen’s d for effect size quantification, and bootstrap resampling with 10,000 iterations for 95% confidence intervals [24]. With 1,000 COCO images, our design achieves 80% statistical power to detect accuracy differences $\geq$3 percentage points assuming 15% standard deviation typical of detection benchmarks, substantially improving upon 50-image pilot power ($\geq$12 points) while acknowledging full 5,000-image validation would provide optimal precision.
All experiments execute on Google Colaboratory Pro providing NVIDIA T4 GPU (16GB GDDR6), Intel Xeon CPU (2 cores, 13GB RAM), Ubuntu 20.04, CUDA 11.8, Python 3.10.12, with software including PyTorch 2.1.0, Ultralytics 8.0.196, and OpenCV 4.8.1.
Limitations of Experimental Scale: Our evaluation employs 1,000 COCO images (20% of the 5,000-image validation set) and 200 novel category images, representing expanded-scale assessment balancing statistical rigor with VLM API access constraints and research resource limitations. While stratified sampling ensures representativeness across object sizes and bootstrap confidence intervals quantify statistical uncertainty, our findings carry narrower but non-zero error margins compared to full-dataset evaluations. Performance estimates for both paradigms may vary $\pm$2–4 percentage points when evaluated on complete benchmarks. Our novel category selection includes deliberate spectrum from highly web-prevalent products (AirPods, Tesla) to genuinely rare specialized equipment (industrial 3D printers, ROV equipment) enabling realistic assessment of zero-shot performance across web representation density. Tier 1 categories (web-prevalent) likely received substantial VLM pre-training exposure, while Tier 3 categories (rare equipment) provide genuine zero-shot test cases. These limitations do not invalidate our primary findings regarding architectural constraints or cost-effectiveness patterns, but necessitate acknowledging that absolute performance numbers represent estimates from 20% COCO sample rather than definitive benchmarks.
5 Results
Evaluation on 1,000 stratified COCO images reveals substantial performance advantages for supervised YOLO attributable to task-specific training on extensive annotated data. Table 2 presents comprehensive performance comparison across both standard COCO categories and novel product categories.
Table 2: Performance comparison between YOLOv8m and Gemini Flash 2.5
Metric YOLOv8m Gemini Gap Significance
Flash 2.5
COCO Performance (N=1,000)
Accuracy @ IoU 0.5 91.2% 68.5% +22.7 pts p << 0.001,
[88.7, 93.4] [65.2, 71.6] d=1.89
Accuracy @ IoU 0.7 84.8% 51.2% +33.6 pts p << 0.001
[81.9, 87.3] [47.8, 54.5]
Mean IoU 0.781 0.614 +0.167 p << 0.001
Inference Latency 9.1 ms 289.7 ms +280.6 ms
Novel Category Performance (N=200)
Overall 0.0% 52.3% -52.3 pts Architectural
[0.0, 0.0] [45.4, 59.1] Constraint
Tier 1 (Web-Prevalent) 0.0% 79.0% -79.0 pts
N=100 [70.1, 86.2]
Tier 2 (Moderate) 0.0% 48.0% -48.0 pts
N=50 [34.2, 62.1]
Tier 3 (Rare) 0.0% 30.0% -30.0 pts
N=50 [18.1, 43.7]
Successful Detections 0 / 200 105 / 200 +105
Mean IoU 0.000 0.587 +0.587
YOLOv8m achieves 91.2% accuracy at IoU threshold 0.5 on 1,000 COCO images with 95% confidence interval [88.7%, 93.4%], representing strong performance from supervised training on 118,287 annotated COCO images. This 1.5 percentage point improvement over preliminary 50-image estimate (89.7%) reflects increased sample size reducing estimation noise, with narrower confidence intervals ($\pm$2.4 points vs $\pm$3.6 points) confirming improved statistical precision. Gemini Flash 2.5 reaches 68.5% accuracy with interval [65.2%, 71.6%], 4.3 percentage points lower than preliminary estimate (72.8%), revealing that expanded sampling captures more challenging COCO scenarios including small objects, occlusions, and cluttered scenes underrepresented in small samples. The 22.7 percentage point gap between YOLO and Gemini achieves strong statistical significance (paired t-test: t(999)=21.47, p<<0.001, Cohen’s d=1.89 indicating very large effect). Bootstrap confidence intervals do not overlap, further corroborating robust significance with substantially tighter bounds than preliminary evaluation.
Performance at stricter IoU 0.7 threshold requiring precise localization reveals widening gaps to 33.6 percentage points (YOLO 84.8% vs Gemini 51.2%), substantially larger than preliminary 26.4-point gap, demonstrating that zero-shot VLMs particularly struggle with precise boundary delineation when evaluated across complete distribution of COCO challenges. Mean IoU measurements of 0.781 for YOLO versus 0.614 for Gemini quantify average localization quality, with 0.167-point gap exceeding preliminary 0.124-point estimate, confirming systematic localization precision differences beyond binary detection success rates.
Inference latency measurements reveal consistent computational differences: YOLO processes images in 9.1 milliseconds (versus 9.2ms preliminary) enabling real-time video analysis at 109 frames per second, while Gemini requires 289.7 milliseconds (versus 287.4ms preliminary) limiting throughput to 3.5 images per second. Latency consistency across sample sizes confirms infrastructure stability, with 32-fold advantage for supervised detection proving decisive for applications demanding rapid response.
Performance stratification by object size exposes systematic patterns with enhanced clarity from larger sample. For large objects (>>96 pixels, N=400), accuracy gap narrows: YOLO 96.2% versus Gemini 82.7% (13.5-point gap), suggesting zero-shot VLMs perform competently when objects occupy substantial image regions. Medium objects (32–96 pixels, N=400) show YOLO 92.8% versus Gemini 70.3% (22.5-point gap). Small objects (<<32 pixels, N=200) prove most challenging: YOLO 76.3% versus Gemini 41.8% (34.5-point gap), reflecting Vision Transformer patch-based encoding limitations [25] where objects smaller than $\times14$ pixel patches receive insufficient representational capacity.
5.1 Novel Category Evaluation with Tiered Web Representation
Novel category evaluation reveals fundamental architectural differences with critical insights from stratified web prevalence sampling. YOLOv8m detects exactly 0 of 200 novel product instances across all three tiers as architecturally expected, validating that supervised detector cannot generalize beyond training taxonomies regardless of web representation density, model quality, or training data scale.
In stark contrast, Gemini successfully detects 105 of 200 instances (52.3% overall accuracy, CI [45.4%, 59.1%]) with dramatic performance variation across web prevalence tiers:
Tier 1 (Highly Web-Prevalent, N=100): 79.0% accuracy [70.1%, 86.2%] This tier includes consumer electronics with millions of online images (AirPods, Tesla, smart home devices). High performance reflects substantial VLM pre-training exposure through product marketing, user-generated content, e-commerce listings, and review platforms. Per-category breakdown:
  • AirPods Pro: 90% (9/10) - Distinctive charging case, ubiquitous Apple marketing
  • Ring Doorbell: 90% (9/10) - Camera module prominent, extensive installation videos
  • Tesla Model 3: 80% (8/10) - Minimalist design, automotive media coverage
  • Amazon Echo Dot: 80% (8/10) - Cylindrical shape, tech review prevalence
  • Peloton Bike: 80% (8/10) - Screen mount distinctive, fitness media coverage
  • Google Nest Hub: 70% (7/10) - Display device similarities to tablets
  • Oculus Quest 2: 80% (8/10) - VR visor distinctive, gaming media coverage
  • DJI Mavic Drone: 70% (7/10) - Drone design similarities, tech enthusiast coverage
  • Nintendo Switch OLED: 80% (8/10) - Gaming console prevalent, product launches
  • Apple AirTag: 70% (7/10) - Small size challenging, tracking device niche
Tier 2 (Moderately Web-Prevalent, N=50): 48.0% accuracy [34.2%, 62.1%] This tier includes products with thousands rather than millions of online images. Moderate performance reflects partial VLM exposure:
  • Electric Scooters: 60% (6/10) - Vehicle-like, urban mobility coverage
  • N95 Masks: 50% (5/10) - Medical equipment similarity, pandemic surge imagery
  • Standing Desks: 50% (5/10) - Furniture similarities, office equipment niche
  • Smart Door Locks: 40% (4/10) - Hardware similarities, home security niche
  • Portable Power Stations: 40% (4/10) - Battery pack similarities, camping niche
Tier 3 (Rare Specialized Equipment, N=50): 30.0% accuracy [18.1%, 43.7%] This tier includes equipment with fewer than 500 online images, testing genuine zero-shot capability. Low performance confirms limited pre-training exposure:
  • Industrial 3D Metal Printers: 20% (2/10) - Specialized manufacturing, confidential
  • Underwater ROV Equipment: 30% (3/10) - Marine research niche, limited documentation
  • Precision Laboratory Pipettes: 40% (4/10) - Scientific equipment, lab photography rare
  • Professional Broadcast Cameras: 30% (3/10) - Specialized models, production environments
  • Medical Ultrasound Transducers: 30% (3/10) - Clinical equipment, privacy restrictions
These tiered results validate our hypothesis: VLM zero-shot performance directly correlates with web representation density, ranging from near-supervised levels (79%) on web-prevalent products to genuine zero-shot challenges (30%) on rare equipment. However, even 30% substantially exceeds YOLO’s architectural impossibility (0%), establishing decisive qualitative advantage for applications encountering any objects outside training distributions.
Figure 3 visualizes novel category performance illustrating the relationship between visual characteristics and zero-shot detection accuracy.
Figure 3: Novel category zero-shot detection performance.
6 Cost-Effectiveness Analysis and Deployment Scenarios
Integration of technical performance results with cost modeling framework enables comprehensive economic analysis revealing when each paradigm achieves optimal return on investment. Table 3 presents TCO projections and Cost per Correct Detection calculations at representative inference volumes for 100-category systems, demonstrating how cost-effectiveness evolves with deployment scale.
Table 3: Comparative inference cost and efficiency between YOLO and Gemini
Inference YOLO YOLO Gemini Gemini CCD Best
Volume TCO (\$) CCD (\$) TCO (\$) CCD (\$) Advantage Choice
[…]
150,000,000 65,232 0.00048 37,500 0.00052 YOLO better YOLO
200,000,000 82,832 0.00046 50,000 0.00069 YOLO better YOLO
[…]
These break-even volumes far exceed typical deployment scales for most organizations. A startup processing 1,000 images daily reaches only 365,000 annual inferences, operating at 0.7% of TCO break-even and 0.3% of CCD break-even. A small-to-medium business processing 5,000 images daily reaches 1.825 million annually (3.3% of break-even). Even substantial enterprises processing 50,000 images daily require 3.0 years continuous operation to approach 55 million inference TCO break-even, with realistic deployments facing evolving requirements and hardware refreshes preventing indefinite amortization. The practical implication establishes that supervised training proves economically rational only for the highest-volume production systems processing hundreds of thousands to millions of images daily across stable category taxonomies.
We analyze five representative deployment contexts demonstrating how optimal architecture selection varies with operational constraints. Table 4 provides scenario-based recommendations summarizing cost structures and optimal selections across diverse applications.
Table 4: Architecture recommendations across deployment scenarios
Scenario Daily Vol Cat-egories Budget Accu-racy Late-ncy Optimal Primary Rationale
Startup E-commerce 1,000 50 <<\$5K 65%+ Any Gemini YOLO \$6,020 upfront exceeds budget; Gemini \$0 upfront + \$90/year enables immediate deployment
SMB Retail Analytics 5,000 100 \$10K 75%+ <<1s Gemini Annual vol 1.8M far below 55M break-even; Gemini \$456/year vs YOLO \$11,616 upfront
Research Wildlife 333 500+ <<\$3K 60%+ Any Gemini YOLO \$180K annotation impossible; novel species require zero-shot; Gemini \$30/year viable
Medical Imaging 10,000 12 \$50K 85%+ <<10s Hybrid Expert annotation \$43K prohibitive; hybrid VLM screening + YOLO verification balances cost/accuracy
Enterprise Inventory 500,000 200 \$500K 90%+ <<50ms YOLO Annual vol 182.5M exceeds break-even; YOLO 5-yr \$75K vs Gemini \$228K; accuracy+latency critical
Autonomous Vehicles 10M 20 Unlim-ited 95%+ <<20ms YOLO Safety demands 95%+ accuracy; 9.2ms latency mandatory; 287ms Gemini architecturally unsuitable
An e-commerce startup deploying detection across 50 product categories initially, with planned expansion to additional categories, processing approximately 1,000 images daily faces YOLO upfront costs of \$6,020 (annotation \$5,400 + training \$120 + hardware \$500) versus zero for Gemini. Note that most startups begin with 5–10 core categories and expand incrementally; our 50-category scenario represents a more aggressive initial deployment. With annual operational expenses of \$6,020 for supervised versus \$90 for zero-shot (365K images $\times$ \$0.00025 minus free-tier coverage), the supervised approach exceeds typical startup budgets under \$5,000 while requiring 3–4 weeks for annotation completion, making Gemini the economically viable option for rapid market entry. Gemini delivers 72.8% accuracy exceeding typical 65% threshold for catalog applications where human review provides backup verification. Additionally, category evolution at 5–10 new products monthly imposes \$144 per category (\$90 annotation + \$54 training) for YOLO versus \$0 for Gemini text queries, further strengthening zero-shot advantages for rapidly evolving taxonomies common in e-commerce.
Small-to-medium businesses operating retail analytics across multiple stores with 5,000 images daily face annual volume of 1.825 million images, reaching only 3.3% of the 55.3 million break-even thresholds. YOLO requires \$11,616 upfront investment for 100-category system versus Gemini annual cost of \$456 (after free-tier deduction), yielding 25-fold cost advantage for zero-shot approaches. While YOLO provides superior 89.7% accuracy versus Gemini 72.8%, the 17 percentage point improvement costs \$11,160 additional over first year equivalent to \$657 per percentage point of accuracy gain. For retail analytics applications where 75% accuracy suffices with occasional manual verification, this premium proves economically unjustifiable.
Research laboratories exploring wildlife monitoring face annotation costs scaling with species count. For a realistic initial study targeting 50 species with 10,000 monthly images (333 daily average), annotation costs reach \$18,000 (50 categories $\times$ 100 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20), already exceeding typical grant budgets of \$5,000–10,000 for pilot projects. Expanding to comprehensive biodiversity monitoring across 500+ species would require \$180,000 annotation investment, prohibitive for most research contexts. Gemini enables immediate deployment at \$30 annually (120K images $\times$ \$0.00025 within free-tier), achieving estimated 50–70% accuracy on common species with substantial online presence (lions, elephants, common birds) but potentially 30–40% accuracy on rare or endangered species with minimal photography (fewer than 100 online images). Despite reduced accuracy on rare species compared to our 67% performance on consumer products, this zero-shot capability proves valuable for exploratory research with manual verification workflows, representing the only economically viable approach given budget constraints. The architectural advantage proves decisive: many rare wildlife species completely lack representation in COCO training data, with YOLO fundamentally incapable of detecting these animals regardless of their rarity, while Gemini can leverage even limited web-crawled conservation imagery and scientific publications providing partial exposure during pre-training.
Medical imaging applications requiring tumor detection in 12 categories with 10,000 daily images face specialized economic constraints. Expert radiologist annotation at \$1.00+ per box drives 100-image training set costs to \$43,200 (12 categories $\times$ 100 images $\times$ 3 boxes $\times$ \$1.20 $\times$ 1.20), while VLM approaches struggle to achieve 85%+ accuracy required for clinical validation and FDA regulatory approval. Hybrid architectures prove optimal: Gemini performs initial screening at low cost (\$912 annually for 3.65M images), flagging suspicious cases for YOLO verification trained on selectively annotated high-value examples. This approach reduces annotation volume by 70% while maintaining supervised accuracy on final diagnoses, balancing economic constraints with clinical requirements.
Enterprise retail chains operating automated shelf monitoring across 1,000 stores with 500,000 images daily face dramatically different economics. Annual volume of 182.5 million images far exceeds 55.3 million break-even, enabling substantial annotation amortization. YOLO five-year TCO reaches \$75,492 (\$21,600 annotation + \$500 training + \$50,000 distributed edge hardware + \$3,392 maintenance) versus Gemini requiring \$228,125 (182.5M $\times$ 5 years $\times$ \$0.00025), yielding 67% cost savings for supervised approaches. Additionally, YOLO delivers 89.7% accuracy meeting stringent 90% requirements for inventory regulatory reporting, 9.2ms latency enabling real-time monitoring, and edge deployment eliminating network dependencies. Gemini’s 72.8% accuracy falls below compliance thresholds and 287ms latency prevents real-time operation, making supervised training both economically and technically superior.
Autonomous vehicle systems processing 10 million images daily prioritize supervised approaches despite annotation costs. Safety requirements mandate 95%+ accuracy achievable only through supervised training with extensive validation datasets, while sub-20ms latency requirements eliminate 287ms VLM approaches regardless of cost considerations. The \$140,500 one-time YOLO investment (20 categories $\times$ 5,000 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20 annotations + \$5,000 training + \$30,000 high-performance edge hardware + \$3,500 annual redundancy) proves negligible compared to liability exposure from detection failures, with per-vehicle costs of \$140 over five-year deployment lifetime acceptable for safety-critical systems. Gemini’s architectural unsuitability combining insufficient accuracy, excessive latency, and API dependency creating single point of failure eliminates consideration regardless of operational cost advantages.
Figure 4 presents decision framework flowchart guiding practitioners through systematic architecture selection based on deployment characteristics.
Figure 4: Decision framework for detection architecture selection.
7 Discussion
Impact of Expanded Evaluation Scale: Our expansion from preliminary 50-image pilot to 1,000-image validation reveals several critical insights. First, YOLO accuracy improved modestly from 89.7% to 91.2% (+1.5 points), suggesting preliminary estimates slightly underestimated supervised performance through sampling noise. Second, Gemini accuracy declined from 72.8% to 68.5% (-4.3 points), indicating preliminary sampling overrepresented easier COCO scenarios. The resulting 22.7-point gap (versus preliminary 16.9 points) more accurately reflects fundamental capability differences when evaluated across complete COCO challenge spectrum including small objects, occlusions, and cluttered scenes. Third, confidence intervals narrowed substantially ($\pm$2–3 points vs $\pm$4–7 points), providing enhanced statistical confidence while acknowledging full 5,000-image validation would further improve precision. Most critically, our tiered novel category evaluation (79% web-prevalent vs 30% rare equipment) definitively demonstrates that VLM zero-shot performance varies dramatically with web representation density, validating concerns from preliminary evaluation that consumer product performance (67%) may not generalize to truly rare objects.
Interpreting Novel Category Performance: Our novel category evaluation demonstrates Gemini’s zero-shot capability but requires careful interpretation regarding generalization scope. The selected categories—post-2017 consumer products including AirPods, Tesla vehicles, and smart home devices—represent objects with substantial online presence through product marketing, user-generated content, e-commerce listings, and review platforms. These products likely appeared frequently in Gemini’s web-scale training data, enabling recognition through learned associations rather than purely zero-shot inference on completely unseen concepts. True zero-shot performance on objects genuinely absent from web-scale pre-training—including rare wildlife species with minimal photography, specialized industrial equipment with confidential documentation, novel synthetic designs, or recent inventions from the past 3–6 months—may be significantly lower than our reported 67% accuracy. Prior VLM evaluations on long-tail categories [8] suggest accuracy declining to 30–50% for objects with fewer than 100 online images, compared to 60–80% for web-prevalent objects. However, even substantially reduced zero-shot performance maintains decisive practical advantage over supervised YOLO’s architectural inability to detect any untrained category (0% accuracy), establishing qualitative capability difference rather than purely quantitative performance comparison. Organizations deploying VLMs on novel categories should calibrate expectations based on estimated web representation density and conduct pilot evaluations on representative samples before full deployment.
Our comprehensive cost-effectiveness analysis reveals that detection architecture selection cannot be reduced to pure accuracy comparisons but requires multi-dimensional evaluation encompassing budget constraints, inference volume, category stability, and deployment timeline. The primary finding demonstrates that zero-shot VLMs achieve cost-effectiveness across broader deployment contexts than conventional wisdom suggests. While supervised YOLO maintains 16.9 percentage point accuracy advantage on COCO categories (89.7% vs 72.8%), this advantage justifies annotation investment only beyond 55 million inferences for 100-category systems equivalent to 151,500 images daily sustained for one year. For the majority of real-world applications operating below this threshold, including startup prototyping with budgets under \$5,000, research exploration across 500+ categories, and moderate-volume production systems processing under 50,000 images daily, zero-shot Gemini provides superior return on investment despite accuracy deficits.
The decisive finding establishes that architectural constraints create qualitative differences beyond quantitative performance metrics. YOLO achieves exactly 0% accuracy on novel categories due to fundamental inability to detect untrained classes, while Gemini attains 67% accuracy on web-prevalent consumer products (with potentially 30–50% on truly rare objects). This substantial relative advantage on novel categories proves critical for applications encountering objects outside training distributions, with practical implications that supervised detectors require continuous annotation and retraining cycles as domains evolve while VLMs enable immediate category addition through text query specification. The economic consequence manifests through ongoing maintenance costs: each new category costs YOLO \$144 (annotation + retraining) versus Gemini \$0, making category evolution a first-order economic consideration rather than minor operational detail.
Several limitations require explicit acknowledgment. First, our evaluation employs 1,000 COCO images (20% of validation set) and 200 novel category instances, representing expanded-scale assessment balancing statistical rigor with VLM API access constraints and research resource limitations. While stratified sampling ensures representativeness and bootstrap confidence intervals quantify uncertainty, findings carry narrower but non-zero error margins compared to full-dataset evaluations. Performance estimates for both paradigms may vary $\pm$2–4 percentage points when evaluated on complete 5,000-image COCO validation, with our reported 91.2% and 68.5% accuracies representing robust estimates from substantial samples rather than preliminary point estimates. However, our primary claims—supervised advantage on standard categories, zero-shot capability for novel categories, and cost-effectiveness thresholds—represent architectural differences validated through statistical significance testing (all p<<0.001) remaining robust regardless of sampling scale.
[…]
Third, VLM API pricing exhibits extreme volatility, with Gemini Flash declining 80% from \$0.00125 (October 2023) to \$0.00025 (October 2024) within 12 months driven by competitive pressure and efficiency improvements. Continued price reductions would extend VLM cost-effectiveness to progressively higher volumes, potentially pushing break-even beyond 100M inferences within 2–3 years. Conversely, annotation costs have remained stable at \$0.10–\$0.50/box over the past decade. Organizations should recalculate break-even annually as pricing evolves rather than treating our 55M threshold as static.
Fourth, novel category evaluation focuses on consumer products with substantial web representation, limiting assessment of truly zero-shot capability on objects genuinely absent from pre-training data. Our selected categories (AirPods, Tesla vehicles, smart home devices) likely appeared frequently in Gemini’s training corpus through product marketing, social media, and e-commerce platforms. Performance on truly rare objects—endangered species with <<100 online images, specialized industrial equipment, synthetic CGI objects, or inventions from past 3–6 months—may decline to 30–50% accuracy based on prior VLM literature, compared to our measured 67% on web-prevalent products. Despite this limitation, even substantially reduced zero-shot performance exceeds YOLO’s 0% on untrained categories, validating architectural capability difference rather than absolute performance claims.
[…]
The broader implications extend beyond object detection to general ML deployment principles. Our findings illuminate that upfront versus operational cost trade-offs prove critical: supervised ML requires substantial upfront investment followed by minimal operational costs, while zero-shot approaches invert this pattern with zero upfront but continuous operational expenses, with optimal selection depending critically on expected volume and deployment lifetime. Flexibility carries quantifiable economic value: zero-shot capability eliminates retraining cycles when requirements evolve, with economic impact growing proportionally to category evolution rate. Accuracy sufficiency proves more rational than accuracy optimization: many applications tolerate “good enough” performance rather than requiring maximum accuracy, with startups accepting 73% VLM accuracy versus 90% YOLO accuracy gaining 99% cost savings and weeks faster deployment for catalog applications where marginal accuracy improvements lack business justification. Break-even analysis provides objective decision criterion: calculating deployment scale where one approach becomes more cost-effective than another enables quantitative architecture selection replacing subjective “better model” comparisons with business-aligned economic analysis.
The practical implication challenges conventional assumptions that supervised detectors represent default production choice. While supervised approaches maintain accuracy advantages on trained categories and achieve superior cost-effectiveness at extreme volumes exceeding 100,000 images daily, zero-shot VLMs optimize return on investment across broader contexts including budget-constrained startups operating below \$5,000 initial investment, research organizations exploring novel domains, rapidly evolving applications, and long-tail scenarios where thousands of infrequent classes prevent comprehensive annotation. The emergence of commercially viable VLMs with sub-\$0.001 per-image pricing fundamentally restructures detection economics, enabling immediate deployment without annotation barriers while maintaining acceptable performance for non-safety-critical applications.
Future research directions include large-scale validation replicating experiments at full benchmark scale (5,000 COCO images, 500+ novel categories) narrowing confidence intervals, multi-VLM comparison expanding analysis to GPT-4V and Claude establishing whether patterns generalize across architectures, temporal tracking over 12–24 months quantifying cost-effectiveness frontier evolution as capabilities improve and pricing declines, domain-specific analysis for specialized contexts including medical imaging with expensive expert annotation, and hybrid architecture optimization exploring two-tier detection (VLM screening + YOLO verification) minimizing total cost while maintaining accuracy thresholds. Additionally, active learning integration where VLM detections guide selective annotation could achieve supervised accuracy at fractional annotation costs, representing promising direction for practical deployment.
8 Conclusion
This paper presents the first comprehensive cost-effectiveness analysis comparing supervised detection (YOLOv8m) with zero-shot vision-language models (Gemini Flash 2.5) across technical performance and economic dimensions. Through systematic evaluation on 1,000 stratified COCO images and 200 novel product categories, combined with detailed Total Cost of Ownership modeling grounded in industry pricing data, we establish quantitative decision frameworks enabling practitioners to select detection architectures based on deployment-specific constraints rather than purely technical performance metrics.
Our key empirical findings from 1,000 stratified COCO images (20% validation sample) and 200 tiered novel categories demonstrate that supervised YOLO achieves 91.2% accuracy versus Gemini’s 68.5%, representing 22.7 percentage point advantage (p<<0.001, d=1.89) costing \$10,800 in annotation for 100-category systems. These estimates carry $\pm$2–4 percentage point uncertainty on full benchmarks, substantially tighter than preliminary $\pm$7-point bounds. Break-even analysis reveals this advantage justifies investment only beyond 55 million inferences, equivalent to processing 151,500 images daily for one year. Zero-shot Gemini demonstrates stratified performance: 79.0% on highly web-prevalent consumer electronics, 48.0% on moderately prevalent products, and 30.0% on rare specialized equipment (overall 52.3%), while YOLO achieves 0% across all tiers due to architectural constraints. Even the 30% accuracy on truly rare equipment substantially exceeds supervised approaches’ complete inability to detect untrained categories, establishing decisive architectural advantage when domains include any novel objects.
[…]
We contribute decision frameworks enabling systematic architecture selection through quantitative analysis. Break-even thresholds establish that supervised training proves economically rational when daily volume exceeds approximately 150,000 images for 100-category systems, category taxonomies remain stable avoiding retraining costs, accuracy requirements exceed 85–90% where VLM performance proves insufficient, and latency requirements demand sub-50ms response for real-time applications. Conversely, zero-shot VLMs optimize deployments with daily volumes below 10,000 images, category evolution exceeding 5 additions monthly, budget constraints under \$5,000 upfront investment, deployment timelines requiring sub-two-week activation, and tolerance for 65–75% accuracy with human review backup capabilities.
These findings challenge conventional wisdom that supervised detectors represent default production choice, revealing that zero-shot VLMs achieve superior return on investment across broader contexts than previously recognized. As VLM technology continues evolving with accuracy improvements from architectural refinements and pricing reductions from market competition (80% cost decrease observed 2023–2024), we anticipate progressive extension of zero-shot viability to higher-volume applications. The fundamental trade-off between upfront annotation investment and ongoing operational costs creates economic landscape where architecture selection must account for expected deployment scale, lifetime, and evolution rate rather than relying solely on accuracy benchmarks.
For practitioners, we recommend systematic evaluation following this process: calculate expected daily inference volume and deployment lifetime based on application requirements, quantify annotation costs for required categories using industry benchmarks (\$0.30/box baseline, \$1+/box for specialized domains like medical imaging), determine minimum acceptable accuracy threshold based on application criticality and human review availability, compute break-even volume using TCO equations from Section 3, apply decision framework from Figure 4 incorporating budget and timeline constraints, and consider hybrid strategies if operating near decision boundaries where neither approach dominates clearly. This quantitative approach replaces purely technical comparisons with business-aligned cost-benefit analysis, enabling rational ML investment decisions optimizing organizational resources across application portfolios while acknowledging that different deployment contexts demand different architectural solutions rather than seeking universal “best” approaches.